Given that p(x) = x is a particular solution to the

differential equation y''+ y = x write the generalized sotlution and check by verifying that the solution satisfies

the equation


1
Expert's answer
2021-09-22T22:55:43-0400

Let us solve the differential equation "y''+ y = x." The characteristic equation "k^2+1=0" has the roots "k_1=i" and "k_2=-i." Taking into account that "p(x) = x" is a particular solution, we conclude that the general solution is "y=C_1\cos x+C_2\sin x+x."

Let us show that the solution satisfies the equation. Since "y'=-C_1\sin x+C_2\cos x+1,\ y''=-C_1\cos x-C_2\sin x," and

"y''+ y=-C_1\cos x-C_2\sin x+C_1\cos x+C_2\sin x+x=x,"

we conclude that "y=C_1\cos x+C_2\sin x+x" is indeed the solution of the equation.